第3章確率
References参考文献
Lopez, Shane, Preety Sidhu. “U.S. Teachers Love Their Lives, but Struggle in the Workplace.” Gallup Wellbeing, 2013. http://www.gallup.com/poll/161516/teachers-love-lives-struggle-workplace.aspx (accessed May 2, 2013).
Shane Lopez、Preety Sidhu。「米国の教師は生活に満足しているが、職場では苦闘している」。Gallup Wellbeing、2013年。http://www.gallup.com/poll/161516/teachers-love-lives-struggle-workplace.aspx(2013年5月2日閲覧)。
Data from Gallup. Available online at www.gallup.com/ (accessed May 2, 2013).
Gallupのデータ。オンラインで閲覧可能:www.gallup.com/(2013年5月2日閲覧)。
Chapter Review章の復習
Two events A and B are independent if the knowledge that one occurred does not affect the chance the other occurs. If two events are not independent, then we say that they are dependent.
二つの事象AとBについて、一方が起こったと知っても他方の起こる可能性が変わらないなら、それらは独立である。二つの事象が独立でなければ、従属であるという。
In sampling with replacement, each member of a population is replaced after it is picked, so that member has the possibility of being chosen more than once, and the events are considered to be independent. In sampling without replacement, each member of a population may be chosen only once, and the events are considered not to be independent. When events do not share outcomes, they are mutually exclusive of each other.
復元抽出では、母集団から選んだ各要素を元に戻すので、同じ要素が2回以上選ばれる可能性があり、事象は独立と見なされる。非復元抽出では、母集団の各要素は1回しか選ばれず、事象は独立でないと見なされる。事象に共通する結果がない場合、それらは互いに排反である。
Formula Review公式の復習
If A and B are independent, P(A AND B) = P(A)P(B), P(A|B) = P(A) and P(B|A) = P(B).
AとBが独立なら、P(A AND B) = P(A)P(B)、P(A|B) = P(A)、P(B|A) = P(B)となる。
If A and B are mutually exclusive, P(A OR B) = P(A) + P(B) and P(A AND B) = 0.
AとBが排反なら、P(A OR B) = P(A) + P(B)、P(A AND B) = 0となる。
Exercises 42–45練習問題 42–45
Exercise 42問題 42
E and F are mutually exclusive events. P(E) = 0.4; P(F) = 0.5. Find P(E∣F).
EとFは排反な事象であり、P(E) = 0.4、P(F) = 0.5である。P(E∣F)を求めよ。
Exercise 43問題 43
J and K are independent events. P(J|K) = 0.3. Find P(J).
JとKは独立な事象であり、P(J|K) = 0.3である。P(J)を求めよ。
Solution解き方
P(J) = 0.3
Exercise 44問題 44
U and V are mutually exclusive events. P(U) = 0.26; P(V) = 0.37. Find:
UとVは排反な事象であり、P(U) = 0.26、P(V) = 0.37である。次を求めよ。
- P(U AND V) =
- P(U|V) =
- P(U OR V) =
Exercise 45問題 45
Q and R are independent events. P(Q) = 0.4 and P(Q AND R) = 0.1. Find P(R).
QとRは独立な事象であり、P(Q) = 0.4、P(Q AND R) = 0.1である。P(R)を求めよ。
Solution解き方
P(Q AND R) = P(Q)P(R)
0.1 = (0.4)P(R)
P(R) = 0.25
Homework宿題
Exercises 46–57練習問題 46–57
Use the following information to answer the next 12 exercises. The graph shown is based on more than 170,000 interviews done by Gallup that took place from January through December in a certain year. The sample consists of employed Americans 18 years of age or older. The Emotional Health Index Scores are the sample space. We randomly sample one Emotional Health Index Score.
英語のヒント
次の情報を使い、続く12の問題に答えよ。このグラフは、ある年の1月から12月にGallupが行った17万件を超える面接調査に基づいている。標本は、就業している18歳以上の米国人からなる。情緒的健康指標の得点を標本空間とする。その得点を一つ無作為に抽出する。
Figure 3.4図 3.4
Exercise 46問題 46
Find the probability that the selected occupation has an Emotional Health Index Score of 82.7.
選ばれた職業の情緒的健康指標の得点が82.7である確率を求めよ。
Exercise 47問題 47
Find the probability that the selected occupation has an Emotional Health Index Score of 81.0.
選ばれた職業の情緒的健康指標の得点が81.0である確率を求めよ。
Exercise 48問題 48
Find the probability that the selected occupation has an Emotional Health Index Score more than 81.
選ばれた職業の情緒的健康指標の得点が81を超える確率を求めよ。
Exercise 49問題 49
Find the probability that the selected occupation has an Emotional Health Index Score between 80.5 and 82.
選ばれた職業の情緒的健康指標の得点が80.5から82の間にある確率を求めよ。
Exercise 50問題 50
If we know an occupation has an Emotional Health Index Score of 81.5 or more, what is the probability that it is 82.7?
ある職業の情緒的健康指標の得点が81.5以上だと分かっているとき、その得点が82.7である確率はいくらか。
Exercise 51問題 51
What is the probability that an occupation has an Emotional Health Index Score of 80.7 or 82.7?
ある職業の情緒的健康指標の得点が80.7または82.7である確率はいくらか。
Exercise 52問題 52
What is the probability that an occupation has an Emotional Health Index Score less than 80.2 given that it is already less than 81?
ある職業の情緒的健康指標の得点が81未満であるという条件のもとで、80.2未満である確率はいくらか。
Exercise 53問題 53
What occupation has the highest Emotional Health Index Score?
情緒的健康指標の得点が最も高い職業は何か。
Exercise 54問題 54
What occupation has the lowest Emotional Health Index Score?
情緒的健康指標の得点が最も低い職業は何か。
Exercise 55問題 55
What is the range of the data?
データの範囲はいくらか。
Exercise 56問題 56
Compute the average EHIS.
情緒的健康指標の得点(EHIS)の平均を計算せよ。
Exercise 57問題 57
If all occupations are equally likely for a certain individual, what is the probability that they will have an occupation with lower than average EHIS?
ある人にとってすべての職業が同様に確からしいなら、EHISが平均より低い職業に就く確率はいくらか。
Bringing It Together総合問題
Exercises 58–61練習問題 58–61
Exercise 58問題 58
A previous year, the weights of the members of the San Francisco 49ers and the Dallas Cowboys were published in the San Jose Mercury News. The factual data are compiled into Table 3.3.
ある過去の年に、サンフランシスコ49ersとダラス・カウボーイズの選手の体重がSan Jose Mercury Newsに掲載された。その実際のデータを表3.3にまとめた。
Shirt# 背番号 |
≤ 210 |
211–250 |
251–290 |
290≤ |
|---|---|---|---|---|
1–33 |
21 |
5 |
0 |
0 |
34–66 |
6 |
18 |
7 |
4 |
66–99 |
6 |
12 |
22 |
5 |
For the following, suppose that you randomly select one player from the 49ers or Cowboys.
以下では、49ersまたはカウボーイズから選手を1人無作為に選ぶものとする。
If having a shirt number from one to 33 and weighing at most 210 pounds were independent events, then what should be true about P(Shirt# 1–33|≤ 210 pounds)?
背番号が1から33である事象と、体重が210ポンド以下である事象が独立なら、P(背番号1–33|210ポンド以下)について何が成り立つはずか。
Exercise 59問題 59
The probability that a person develops cancer is 0.4567. The probability that a person has at least one false positive test result (meaning the test comes back for cancer when the person does not have it) is 0.51. Some of the following questions do not have enough information for you to answer them. Write “not enough information” for those answers. Let C = a person develops cancer and P = a person has at least one false positive.
英語のヒント
ある人ががんを発症する確率は0.4567である。少なくとも1回の偽陽性の検査結果が出る確率は0.51である(偽陽性とは、がんでないのに、検査でがんと判定されることである)。以下の問いの中には、答えるための情報が足りないものがある。それらには「情報が足りない」と答えよ。C = 人ががんを発症する、P = 人が少なくとも1回偽陽性となる、とする。
- P(C) = ______
- P(P|C) = ______
- P(P|C') = ______
- If a test comes up positive, based upon numerical values, can you assume that a male has cancer? Justify numerically and explain why or why not.検査結果が陽性だった場合、数値に基づいて、その男性はがんだと仮定できるか。数値で根拠を示し、その理由を説明せよ。
Solution解き方
- P(C) = 0.4567
- not enough information情報が足りない
- not enough information情報が足りない
- No, because over half (0.51) of men have at least one false positive textいいえ。男性の半数を超える割合(0.51)が、少なくとも1回偽陽性の検査結果を経験するからである。
Exercise 60問題 60
Given events G and H: P(G) = 0.43; P(H) = 0.26; P(H AND G) = 0.14
事象GとHについて、P(G) = 0.43、P(H) = 0.26、P(H AND G) = 0.14が与えられている。
- Find P(H OR G).P(H OR G)を求めよ。
- Find the probability of the complement of event (H AND G).事象(H AND G)の余事象の確率を求めよ。
- Find the probability of the complement of event (H OR G).事象(H OR G)の余事象の確率を求めよ。
Exercise 61問題 61
Given events J and K: P(J) = 0.18; P(K) = 0.37; P(J OR K) = 0.45
事象JとKについて、P(J) = 0.18、P(K) = 0.37、P(J OR K) = 0.45が与えられている。
- Find P(J AND K).P(J AND K)を求めよ。
- Find the probability of the complement of event (J AND K).事象(J AND K)の余事象の確率を求めよ。
- Find the probability of the complement of event (J OR K).事象(J OR K)の余事象の確率を求めよ。
Solution解き方
- P(J OR K) = P(J) + P(K) − P(J AND K); 0.45 = 0.18 + 0.37 - P(J AND K); solve to find P(J AND K) = 0.10P(J OR K) = P(J) + P(K) − P(J AND K)より、0.45 = 0.18 + 0.37 − P(J AND K)。これを解くと、P(J AND K) = 0.10となる。
- P(NOT (J AND K)) = 1 - P(J AND K) = 1 - 0.10 = 0.90P(NOT (J AND K)) = 1 − P(J AND K) = 1 − 0.10 = 0.90。
- P(NOT (J OR K)) = 1 - P(J OR K) = 1 - 0.45 = 0.55P(NOT (J OR K)) = 1 − P(J OR K) = 1 − 0.45 = 0.55。
Dependent Events
従属な事象
If two events are NOT independent, then we say that they are dependent.
二つの事象が独立でない場合、それらは従属であるという。
Conditional Probability of One Event Given Another Event
別の事象が起こったという条件のもとでの、ある事象の条件付き確率
P(A|B) is the probability that event A will occur given that the event B has already occurred.
P(A|B)は、事象Bがすでに起こったという条件のもとで、事象Aが起こる確率である。
